### v1

# Compact: Common-token Optimized Model Pruning Across Channels and Tokens

Abstract
Making LLMs more efficient in memory, latency, and serving cost is crucial for edge deployment, interactive applications, and sustainable inference at scale. Pruning is a key technique toward this goal. However, prior pruning methods are limited: width pruning often breaks the standard transformer layout or requires custom inference code, while depth pruning removes entire layers and can cause abrupt accuracy drops. In this work, we propose Compact, which jointly (i) prunes rare vocabulary to shrink embedding/unembedding and (ii) prunes FFN intermediate channels using common-token–weighted activations, aligning importance with the post-pruning token distribution. Compact enjoys merits of both depth and width pruning, such as: deployment-friendliness (keeps a standard transformer architecture), scale-adaptivity (trade off vocab vs. FFN pruning), training-free operation with competitive pruning time, and strong memory savings alongside throughput gains. Experiments across Qwen, LLaMA, and Gemma families (0.5B–70B) show state-of-the-art downstream task performance at similar or higher pruning ratios, with substantial reductions in parameters, GPU memory, and end-to-end latency.
1 Introduction
Large language models (LLMs) have achieved remarkable performance across a wide range of natural language tasks, but their ever-growing parameter counts—reaching billions to hundreds of billions—make deployment expensive in terms of memory, inference time, and energy cost. To broaden access and enable real-world applications such as on-device inference, classroom use, or latency-sensitive systems, it is crucial to compress LLMs while retaining as much performance as possible.
Pruning has been a major line of compression work, mainly categorized into depth pruning and width pruning. Depth pruning often removes entire transformer blocks (Kim et al., 2024; Song et al., 2024; Gromov et al., 2025), while width pruning trims hidden dimensions such as FFN channels or attention heads (Ma et al., 2023; Ashkboos et al., 2024; An et al., 2024). However, these approaches are limited in two ways: (i) They prune largely mechanistically, without analyzing where parameters are concentrated within LLMs (embeddings, FFNs, or attention). This blind pruning means that methods that work for large LLMs often fail for small language models (SLMs), where parameter distributions differ. (ii) They ignore the linguistic nature of NLP models: not all tokens are equally important, yet pruning typically treats all tokens as if they contribute equally. These oversights lead to non-robust pruning performance, especially across scales.
We begin by systematically analyzing parameter distributions across model families and scales. This reveals a clear pattern: embeddings (vocabulary and unembedding layers) dominate in smaller models, while FFNs dominate in larger models. This explains why prior pruning methods lack robustness across scales—they prune the same way regardless of where redundancy actually lies. A second insight comes from the statistics of natural language: token frequencies follow a Zipfian distribution, meaning that rare tokens occur extremely infrequently and contribute little to downstream performance. Removing such rare tokens from the vocabulary reduces embedding size without significantly affecting accuracy, because language tasks are overwhelmingly driven by common tokens.
These two insights naturally lead to our new approach Compact, a simple but powerful pruning framework with two modules: (i) Vocabulary pruning removes rare tokens and shrinks embedding/unembedding matrices, directly reducing parameters and memory usage, especially in SLMs. (ii) Common-token–weighted FFN pruning further reduces redundancy by scoring channels using activations, but weighting only the common tokens that remain valid after vocabulary pruning. Together, these two complementary modules address the limitations of prior work: pruning is now guided by parameter distribution, respects the linguistic structure of language tasks, and remains compatible with the standard transformer architecture.
We evaluate Compact on diverse LLM families (Qwen 2.5, LLaMA 3/3.1/3.2, and Gemma 3) and across scales from 0.5B to 70B parameters. We test on seven downstream benchmarks (MMLU (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2020), ARC-C/E (Clark et al., 2018; 2018), PIQA (Bisk et al., 2020), GSM8K (Cobbe et al., 2021)) and also measure pruning time, inference throughput, and GPU memory usage. Our experiments highlight three phenomena: (i) Scale robustness: Compact maintains state-of-the-art performance at high pruning ratios even for small models, avoiding the collapse seen in prior work. (ii) Smooth degradation: Unlike depth pruning, which shows abrupt performance drops, Compact degrades gracefully with higher pruning. (iii) End-to-end efficiency: Compact’s pruning yields substantial GPU memory savings and improved throughput.
Our contributions are threefold: i) We provide a systematic analysis of parameter distribution across embeddings, FFNs, and attention, revealing scale-dependent redundancy that prior pruning methods overlook. ii) We propose Compact, a novel pruning method which is linguistically grounded, scale-adaptive, and structure-agnostic. iii) We demonstrate state-of-the-art pruning results across LLM families and scales, showing superior retention on downstream tasks together with clear gains in pruning time, inference efficiency, and GPU memory usage.
2 Related Work
Depth Pruning.
One well-explored branch of structured pruning is depth pruning (Kim et al., 2024; Song et al., 2024; He et al., 2024; Lu et al., 2024; Gromov et al., 2025; Chen et al., 2025), which removes entire transformer blocks from the model. Depth pruning is desirable because it only reduces the number of layers in an LLM while still maintaining a standard transformer architecture, allowing the pruned model to be compatible with a wide variety of existing inference frameworks. Shortened LLaMA (Kim et al., 2024) prunes the layers that minimize the increase in perplexity. SLEB (Song et al., 2024) uses a similar metric, but recalculates each layer’s importance after iteratively pruning one layer at a time. Gromov et al. (2025) prunes layers using an angular similarity metric. LLM-Streamline (Chen et al., 2025) has found that training a lightweight network to replace pruned parameters can mitigate performance loss, significantly outperforming existing methods. However, LLM-Streamline requires training the model to achieve performance improvements, which can take hours or days depending on the model size, as well as significant GPU resources. Hence, in this paper we will focus on training-free methods, which can be completed in minutes on a single GPU instead.
While depth pruning excels in its practicality, the coarse-grained removal of entire layers can lead to significant performance drops. In contrast, Compact removes only specific rows and columns of a weight matrix, offering a more fine-grained approach that improves performance.
Width Pruning.
An alternative approach to structured pruning is width pruning (Ma et al., 2023; Ashkboos et al., 2024; An et al., 2024; Xia et al., 2024; Dery et al., 2024; Gao et al., 2024b; Guo et al., 2025), which prunes each transformer layer in the LLM. There are a variety of methods to prune a transformer layer. SliceGPT (Ashkboos et al., 2024) prunes the hidden size by performing orthogonal matrix transformations and low-rank decomposition. LLM-Pruner (Ma et al., 2023) and LoRAPrune (Zhang et al., 2024) identifies and removes coupled structures using gradient information. FLAP (An et al., 2024) prunes FFN channels and self-attention heads based on a channel stability metric. Bonsai (Dery et al., 2024) builds a linear perturbation model of module importances, at the cost of significantly increased pruning time.
Recent works (Yin et al., 2024; Wang et al., 2025) have shown that non-uniform sparsity can improve pruning performance. However, these width pruning methods introduce structural irregularities that deviate from the transformer architecture. This makes it impractical to deploy width-pruned models, as they require specialized code for inference and are not compatible with standard inference libraries. Compact fixes these issues by ensuring that pruned models maintain a standard transformer architecture, enabling a deployment-friendly width-pruning method that outperforms depth pruning techniques.
3 Proposed Method: Compact
3.1 Analyzing Parameter Distribution in LLMs: Vocabulary vs. FFN vs. Attention
Before designing an effective pruning strategy, we first analyze where parameters are concentrated within modern decoder-only transformers. These models consist of three major groups of parameters: (i) vocab parameters, located in the embedding and unembedding layers, which map between the vocabulary space of size $V$ and the hidden dimension $D$; (ii) attention parameters, from the self-attention blocks; and (iii) FFN parameters, from the feed-forward blocks within each layer.
Formally, the embedding and unembedding layers each contain $VD$ parameters, giving
$N_{\text{vocab}}=2VD,$ (1)
(or $VD$ if tied embeddings are used). Each FFN block contains three projection matrices of size $D\times I$, where $I$ is the intermediate dimension, yielding
$N_{\text{FFN}}=3LDI,$ (2)
for $L$ layers. For attention, the number of parameters depends on whether grouped query attention (GQA) is used (Ainslie et al., 2023). With $H$ denoting the ratio of attention heads to KV heads, the count is
$N_{\text{attention}}=2LD^{2}\left(1+\frac{1}{H}\right).$ (3)
When $H=1$, this reduces to $N_{\text{attention}}=4LD^{2}$.
Figure 1: Parameter Distribution across Qwen 2.5 Models of Different Scales
0.5B1.5B3B7B14B32B72B00.20.20.40.40.60.60.80.811ProportionVocabFFNAttention
Asymptotically, $N_{\text{FFN}}$ and $N_{\text{attention}}$ scale as $O(LD^{2})$, while $N_{\text{vocab}}$ scales only as $O(D)$, as $V$ is kept constant when scaling. Thus, as model size grows, vocab parameters become proportionally smaller. Conversely, for smaller models, vocab parameters can constitute a significant fraction of the total, making large vocabularies suboptimal (Tao et al., 2024). We can observe this effect empirically by calculating the relative proportions of each parameter group on popular model families. Figure 1 shows our empirical analysis on the Qwen 2.5 model family (Yang et al., 2024a), which validates our theoretical analysis. Proportions of other model families can be found in Appendix A.1. This motivates our strategy: vocabulary pruning is an efficient way to reduce parameters, especially in small-to-medium LLMs, while FFN pruning is critical for large models.
3.2 From Rare to Common: Rationale of Vocabulary Pruning
Byte-Pair Encoding (BPE) tokenizers follow Zipf’s law (Zhemchuzhina et al., 2022), where most tokens appear extremely rarely. Since BPE builds its vocabulary by merging frequent token pairs, the rarest tokens naturally appear at the end of the vocabulary list.
We define the set $S$ as the $V-V^{\prime}$ rarest tokens in the vocabulary. While these tokens can be directly removed—by pruning the corresponding rows in the embedding and unembedding matrices and deleting their merge rules—the key insight is conceptual: these tokens will never be generated in the pruned model. Consequently, pruning shifts the effective distribution of inputs toward common tokens. This means that subsequent optimization steps should focus on preserving performance under the common-token distribution rather than the full distribution.
This step is highly efficient: it requires no calibration data or forward passes and introduces no additional hyperparameters.
3.3 Intermediate Pruning under the Common-Token Distribution
Pruning vocabulary parameters alone reduces the embedding size but does not address redundancies in the FFNs, which dominate parameter count in large models. To prune FFNs, we adopt an activation-based criterion.
The standard act2 method (Muralidharan et al., 2024; Sandri et al., 2025) defines the importance of channel $k$ as
$I_{k}=\sum_{i=1}^{N}\bigl{(}\sigma(X_{i}W_{1})\odot X_{i}W_{2}\bigr{)}_{k}^{2},$ (4)
summing squared activations over a calibration dataset. However, this equally weights all tokens, including those in $S$.
Since tokens in $S$ will never appear in the input after pruning, their activations should not guide channel importance. We therefore introduce common act2, a weighted variant:
$I_{k}=\sum_{i=1}^{N}w_{i}\bigl{(}\sigma(X_{i}W_{1})\odot X_{i}W_{2}\bigr{)}_{k}^{2},\quad w_{i}=\begin{cases}0&x_{i}\in S,\\ 1&\text{otherwise}.\end{cases}$ (5)
This ensures that FFN pruning is explicitly optimized for the distribution of common tokens that remain valid in the pruned model.
Table 1: Advantages of Compact.
Depth pruning Width pruning Compact
ShortGPT LaCo SliceGPT 2SSP (ours)
Maintains Arch. $\checkmark$ $\checkmark$ $\checkmark$
Scale-adaptive $\checkmark$ $\checkmark$
Inference speedups $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Arch. agnostic $\checkmark$
Linguistically grounded $\checkmark$
3.4 Compact: Joint Pruning Pipeline
Our proposed method, Compact, integrates vocabulary pruning with common act2-based FFN pruning. Importantly, embedding pruning and channel pruning are not performed sequentially in isolation: knowledge of $S$ (rarest tokens) is first identified, then used to guide intermediate pruning, and finally both vocab and FFN parameters are removed. The full pipeline is given in Algorithm 1.
Algorithm 1 Compact
1:Model $M$, calibration dataset $D$, target vocabulary size $V^{\prime}$, target intermediate size $I^{\prime}$
2:Identify $S\leftarrow$ set of $V-V^{\prime}$ rarest tokens in vocabulary.
3:Run forward passes of $M$ on $D$, collect squared activations.
4:For each channel $k$, compute importance $I_{k}$ using common act2 (Eq. 5).
5:for each layer do
6:  Prune $I-I^{\prime}$ least important channels (remove rows of $W_{1},W_{2}$ and columns of $W_{\text{down}}$).
7:end for
8:Prune vocab parameters: remove final $V-V^{\prime}$ rows of embedding and unembedding matrices; delete merges for tokens in $S$.
9:return pruned model $M^{\prime}$.
Advantages of Compact.
i) Compact is scale-adaptive. By exposing two orthogonal knobs: (i) vocabulary pruning at the embedding/unembedding layers and (ii) common-token–weighted pruning of FFN intermediate channels, Compact can be tuned for small LLMs (emphasize vocab pruning) or large LLMs (emphasize intermediate pruning), or any mix to meet a target budget. This tunability preserves capacity on frequent tokens while enabling strong compression across model scales, all without altering the standard decoder-only architecture. ii) Compact is compatible with LLM frameworks. One weakness of most width pruning methods is that they do not maintain a standard transformer architecture. This is indeed the case with SliceGPT, which prunes the hidden size in all layers except for the final one. As a result, models pruned with SliceGPT are not compatible with the Huggingface transformers library, vLLM, or any other inference engines, limiting its practicality. In this aspect, Compact is more akin to depth pruning methods, since pruning only the vocabulary and intermediate size does not affect the transformer architecture. As a result, Compact models are compatible with all inference engines, making it a practical approach to width pruning. The full advantages of Compact is summarized in Table 1.
4 Experiments
4.1 Experimental Setup
Baselines.
We compare with representative and state-of-the-art i) width pruning methods: SliceGPT, (Ashkboos et al., 2024), 2SSP (Sandri et al., 2025); ii) depth pruning methods: ShortGPT (Men et al., 2024), LaCo (Yang et al., 2024b).
LLMs to prune.
We evaluate on a diverse set of LLMs spanning architectures and scales: (i) Small LLMs: Qwen 2.5–0.5B, LLaMA 3.2–1B, and Gemma 3–1B. This mix covers three distinct families, enabling a robustness assessment across architectures; moreover, small LLMs are particularly challenging to prune, as they are often trained beyond the Chinchilla‐optimal compute–data balance, leaving limited redundancy. Nevertheless, pruning small LLMs is highly valuable for edge/on-device use and in privacy- or bandwidth-constrained settings (healthcare, classrooms, federated clients): it shrinks memory/storage, improves cache locality and end-to-end latency, lowers energy and thermal load, and reduces serving cost and carbon footprint. (ii) Large LLMs: LLaMA 3.1–8B and LLaMA 3.1–70B. Together with the 1B variant above, this suite evaluates pruning effectiveness across models differing by orders of magnitude in scale.
Evaluation tasks.
Following SliceGPT Ashkboos et al. (2024), we evaluate all pruning methods using 5 commonly-used natural language benchmarks: HellaSwag (HeSw) (Zellers et al., 2019), WinoGrande (WiGr) (Sakaguchi et al., 2020), ARC-C (Clark et al., 2018), ARC-E (Clark et al., 2018), and PIQA (Bisk et al., 2020). Since Jaiswal et al. (2024) shows that pruned LLMs have significantly degraded performance on more complex tasks, we add MMLU (Hendrycks et al., 2021) to test the models on more difficult general knowledge tasks, and GSM8K (Cobbe et al., 2021) to assess their performance on generation tasks. This collection enables us to obtain a more comprehensive overview of the model’s performance.
Evaluation Criteria.
Details about the evaluation setup and pruning hyperparameters can be found in Appendix A.2 and Appendix A.3, respectively.
  • We report the number of parameters (Para (M)) of each model after pruning, the percentage of parameters that were removed (Ratio (%)), the mean score of the 7 benchmarks (Avg), and the relative mean score compared to the unpruned dense model (Avg%). Note that SliceGPT does not report post-pruning parameter counts; we therefore report only the stated sparsity.
  • It is standard to evaluate pruned models on perplexity and downstream tasks. However, because we reduce vocabulary size, our perplexity naturally decreases, making comparisons to baselines unfair. Thus, we only report i) performance on downstream tasks, ii) efficiency regarding pruning time, inference time, and memory usage.
4.2 Performance on downstream tasks
4.2.1 Results on smaller LLMs
The results are summarized in Table 2, which supports the following observations:
Compact supports a wide variety of model architectures out-of-the-box.
The official implementations of our baseline pruning approaches support just a few model architectures. For instance, SliceGPT supported LLaMA, OPT, and Phi, LaCo supported LLaMA 2 and Baichuan, and ShortGPT only supported LLaMA. We needed to modify these implementations with architecture-specific changes in order to add support for more modern model families like LLaMA 3 and Qwen 2.5. This is because these architectures can differ significantly in how they handle self-attention, positional embeddings, layer normalization, and other details. The strength of Compact is that it only prunes the vocabulary embeddings and FFN blocks. These two model components have been standardized and remain unchanged across the vast majority of model architectures. As a consequence, Compact is architecture-agnostic and runs out of the box across diverse model families, requiring no model-specific modifications. This is most evident with the Gemma 3 family, which incorporates QK-norm and alternates local sliding-window and global attention layers—optimizations that prevented us from adapting baseline methods. Accordingly, baseline results for Gemma 3 are omitted in Table 2 (third block). In contrast, Compact operates on Gemma 3 without any changes.
Compact outperforms baselines with large margins.
Although prior work reports strong results on larger LLMs, the baseline methods perform poorly on smaller LLMs. On Qwen 2.5–0.5B, GSM8K accuracy collapses for all baselines at only a 10% pruning ratio. Likewise, at 10% pruning, MMLU drops to near-random for all models and baselines, with the sole exception of LaCo on Qwen 2.5–0.5B. Because MMLU and GSM8K are the most demanding tasks in our suite, these trends indicate that existing methods fail to preserve performance on challenging benchmarks for small LLMs. In contrast, Compact delays this collapse: it remains marginally above random on MMLU and GSM8K even at 35% pruning across all models. Across models and pruning ratios, Compact attains the highest mean score and leads on nearly all individual benchmarks, despite using a slightly higher pruning ratio than the baselines. Notably, at 35% pruning, Compact maintains similar performance to the baselines at 20%, demonstrating superior robustness under high compression.
Table 2: Pruning smaller LLMs (Qwen 2.5-0.5B, LLaMA 3.2–1B, and Gemma 3–1B) at a $\sim$10%, $\sim$20%, and $\sim$35% ratio. Please note pruning baselines cannot be applied to Gemma 3
Method Para (M) Ratio (%) MMLU HeSw WiGr ARC-C ARC-E PIQA GSM8K Avg Avg% Qwen 2.5–0.5B Dense 494 0.00 47.3 52.2 56.4 32.3 58.2 69.9 34.9 50.2 100.0 Random - - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 57.0 ShortGPT 449 9.11 27.8 44.0 53.0 25.9 45.8 66.8 0.2 37.6 75.1 LaCo 449 9.11 46.1 45.5 56.3 28.2 51.6 65.5 0.4 41.9 83.6 SliceGPT - 10.71 23.2 43.2 53.3 26.4 50.3 63.8 0.0 37.2 74.1 2SSP 444 10.12 25.5 46.6 54.7 27.8 52.2 68.7 1.9 39.6 79.0 Compact 439 11.13 45.2 51.9 55.3 32.4 59.5 70.1 28.7 49.0 97.7 ShortGPT 405 18.02 25.0 37.7 52.0 27.1 41.7 62.1 0.0 35.1 70.0 LaCo 405 18.02 24.0 36.3 49.9 23.5 41.7 62.9 0.0 34.0 67.8 SliceGPT - 19.64 23.1 32.1 52.0 20.2 33.4 53.7 0.0 30.6 61.1 2SSP 397 19.64 24.3 40.9 53.8 25.3 43.9 64.3 0.5 36.1 72.0 Compact 394 20.24 44.1 48.1 55.4 30.6 53.3 66.6 26.3 46.3 92.4 ShortGPT 315 36.23 24.3 27.8 50.1 25.7 26.2 51.8 0.0 29.4 58.7 LaCo 315 36.23 23.9 28.2 47.9 23.9 30.6 56.2 0.0 30.1 60.0 SliceGPT - 36.61 23.1 29.0 51.5 22.5 30.7 53.4 0.0 30.0 59.9 2SSP 315 36.23 22.9 31.3 49.6 22.9 33.8 59.0 0.0 31.4 62.5 Compact 311 37.04 25.5 40.0 53.8 25.2 40.0 62.2 0.5 35.3 70.4 LLaMA 3.2–1B Dense 1236 0.00 36.6 63.8 60.7 36.2 60.7 74.5 5.4 48.3 100.0 Random - - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 59.2 ShortGPT 1112 10.03 23.4 48.0 60.2 30.4 49.8 68.0 0.0 40.0 82.8 LaCo 1112 10.03 24.4 48.4 52.0 29.5 47.9 69.3 0.5 38.9 80.5 SliceGPT - 10.16 23.1 49.4 54.0 28.6 43.4 64.0 0.0 37.5 77.7 2SSP 1114 9.87 30.5 55.5 57.9 32.9 56.7 72.9 3.0 44.2 91.6 Compact 1112 10.03 36.7 61.1 59.7 35.1 57.1 71.9 6.1 46.8 97.0 ShortGPT 993 19.66 22.8 40.0 55.3 29.9 35.2 58.9 0.0 34.6 71.7 LaCo 993 19.66 23.0 35.7 52.4 25.9 37.0 62.4 0.4 33.9 70.1 SliceGPT - 20.31 23.0 39.9 52.3 26.2 38.9 58.6 0.0 34.1 70.7 2SSP 993 19.66 26.8 46.9 53.9 27.1 50.4 68.1 2.2 39.3 81.5 Compact 989 19.98 30.6 54.4 58.6 32.0 51.3 69.9 3.1 42.8 88.8 ShortGPT 810 34.47 24.3 32.5 50.0 28.4 28.9 55.4 0.0 31.4 65.0 LaCo 810 34.47 23.2 37.3 50.1 23.9 29.3 55.3 0.0 29.9 61.9 SliceGPT - 35.16 23.0 30.4 51.2 22.0 32.9 53.4 0.0 30.4 63.1 2SSP 808 34.63 22.9 35.1 52.6 24.5 38.9 60.9 0.0 33.6 69.6 Compact 803 35.03 27.9 42.8 55.6 27.7 41.7 60.6 1.8 36.9 76.4 Gemma3–1B Dense 1009 0.00 24.9 62.1 59.0 38.2 71.9 74.8 2.4 47.6 100.0 Random - - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 60.0 Compact 908 10.01 24.9 60.3 59.0 39.1 69.0 74.1 1.7 46.9 98.4 807 20.02 25.0 55.4 59.0 37.9 63.3 70.0 1.7 44.6 93.6 656 34.99 24.2 45.1 55.9 26.5 46.4 65.3 0.5 37.7 79.2
4.2.2 Results on larger LLMs
LaCo failed to prune LLaMA 3.1-70B due to OOM errors, so it is omitted from our results. Our results are summarized in Table 3.
Compact is robust across scales via complementary pruning.
We see that Compact achieves state-of-the-art performance for larger models as well, indicating that our method is highly robust to a wide range of model sizes. We attribute this robustness to our dual approach to width pruning. As model size increases, the proportion of vocabulary parameters decreases, which decreases the effectiveness of pruning vocabulary size. However, intermediate pruning becomes more effective as model size increases, similarly to most pruning methods (Xu et al., 2024). These two techniques complement each other’s strengths, leading to a robust hybrid pruning method. Although the performance gap between Compact and 2SSP narrows at the 70B size, we note that 2SSP’s hybrid approach of pruning FFN channels and entire attention blocks deviates from a transformer architecture, while Compact’s approach does not. This makes Compact more practical to deploy, while also achieving slightly higher performance.
Compact shows smooth, spike-free degradation.
On LLaMA 3.1–8B, we observe an anomaly on MMLU: ShortGPT surpasses Compact at a 20% pruning ratio, but not at 10% or 35%, with its 20% score even exceeding its 10% score. This aligns with the step-like behavior reported by Gromov et al. (2025) for depth-pruning methods, wherein performance remains nearly intact up to a critical pruning threshold and then collapses abruptly. ShortGPT appears to exhibit this pattern, explaining the spike at 20%. In contrast, Compact (a width-pruning method) shows smooth MMLU degradation as pruning increases—hence the dip relative to ShortGPT at 20%, followed by recovery and a lead at 35%. A similar effect is seen on LLaMA 3.1–70B, where ShortGPT again outperforms on MMLU. Other benchmarks do not display step-like behavior, instead showing either smooth declines or immediate collapses. Notably, even at 35% pruning, Compact remains above random on both MMLU and GSM8K.
Table 3: Pruning larger LLMs (LLaMA 3.1–8B & LLaMA 3.1–70B) at a $\sim$10%, $\sim$20%, and $\sim$35% ratio.
Method Para (M) Ratio (%) MMLU HeSw WiGr ARC-C ARC-E PIQA GSM8K Avg Avg% LLaMA 3.1–8B Dense 8030 0.00 63.4 78.9 73.6 53.4 80.9 81.1 51.6 69.0 100.0 Random - - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 41.4 ShortGPT 7158 10.86 58.0 73.8 70.2 47.4 71.2 77.5 29.3 61.1 88.5 LaCo 7158 10.86 58.8 73.3 72.3 48.9 73.7 76.3 32.0 62.2 90.1 SliceGPT - 10.16 43.9 65.6 67.2 39.4 66.2 71.0 0.0 50.5 73.2 2SSP 7158 10.86 54.1 74.6 71.6 46.8 71.6 79.7 14.1 58.9 85.4 Compact 7227 10.00 59.6 75.2 73.7 50.3 74.9 78.4 27.8 62.9 91.1 ShortGPT 6503 19.02 58.6 64.9 68.4 42.2 58.3 71.6 0.6 52.1 75.5 LaCo 6503 19.02 24.1 54.0 55.3 29.2 51.1 72.4 0.4 40.9 59.3 SliceGPT - 20.12 24.5 51.4 61.8 30.3 49.0 61.9 0.0 39.8 57.8 2SSP 6425 19.99 37.4 67.2 68.4 38.1 61.8 76.8 4.3 50.6 73.3 Compact 6424 20.00 50.7 69.9 70.1 42.8 66.0 75.9 10.8 55.2 80.0 ShortGPT 5195 35.31 23.2 34.3 59.1 29.7 36.9 57.2 0.0 34.4 49.8 LaCo 5195 35.31 23.1 34.8 53.1 27.3 31.5 58.8 0.0 32.7 47.3 SliceGPT - 35.16 23.0 35.0 54.3 23.5 37.2 55.0 0.0 32.6 47.2 2SSP 5238 34.77 25.3 49.9 59.3 27.1 44.3 68.7 2.3 39.5 57.3 Compact 5220 34.99 35.9 56.0 63.3 30.8 48.4 70.6 1.7 43.8 63.5 LLaMA 3.1–70B Dense 70.6 0.00 75.2 85.0 79.5 64.7 86.7 84.4 80.6 79.4 100.0 Random - - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 36.0 ShortGPT 63.7 9.77 75.0 82.9 78.5 60.8 84.8 83.4 74.5 77.1 97.1 SliceGPT - 10.06 70.6 75.4 76.6 58.3 82.0 79.7 0.0 63.2 79.6 2SSP 63.6 9.92 73.4 84.7 78.6 63.9 85.4 84.2 75.1 77.9 98.1 Compact 63.5 10.06 73.5 84.5 79.5 64.0 86.0 84.1 76.0 78.2 98.5 ShortGPT 56.0 20.68 74.9 79.3 78.0 56.0 80.1 80.1 53.0 71.6 90.2 SliceGPT - 20.02 63.1 64.8 74.0 52.4 76.7 73.7 0.0 57.8 72.8 2SSP 56.4 20.11 68.8 83.7 76.2 59.5 82.1 83.7 62.1 73.7 92.8 Compact 56.4 20.11 70.6 83.3 76.2 59.7 82.6 83.7 62.6 74.1 93.3 ShortGPT 44.9 36.40 71.2 66.5 75.7 46.8 69.6 72.3 0.0 57.4 72.3 SliceGPT - 35.06 29.3 37.1 66.6 34.5 59.3 62.6 0.0 41.3 52.0 2SSP 45.8 35.13 58.5 77.7 72.1 50.7 74.2 81.6 25.0 62.8 79.1 Compact 45.9 34.99 59.6 78.1 72.8 53.2 76.1 81.3 25.0 63.7 80.2
4.3 Efficiency in Pruning Time, Inference Latency, and Memory
Table 4: Pruning time ($m$:$s$) comparison at a 35% pruning ratio of LLaMA 3.1.
Method Pru. Time
8B ShortGPT 0:18
LaCo 0:05
SliceGPT 10:48
2SSP 1:26
Compact 0:32
70B ShortGPT 2:10
LaCo OOM
SliceGPT 84:38
2SSP 13:48
Compact 2:17
Pruning time.
The main strength of training-free pruning methods is that they are able to prune large models in a short amount of time. To measure this, we report how long our method takes to prune models at a 35% pruning ratio compared to our baselines. We report our results on LLaMA 3.1-8B and 70B, as larger models enable us to better discriminate between different methods. For fairness, we exclude the time taken to load and save the models and data. All reported results are an average of 3 different pruning runs.
Our results are in Table 4. With LLaMA 3.1-8B, Compact is 3 times faster than 2SSP, our strongest baseline, and within an order of magnitude of our depth pruning methods, pruning the model in less than a minute. At the 70B size, Compact now becomes 6 times faster than 2SSP, and the gap between ShortGPT and Compact shrinks to under 10 seconds. The low pruning times show that Compact has comparable efficiency to our baselines.
Inference speed and memory usage.
Table 5: Throughput and memory usage for LLaMA 3.1-8B.
Method Ratio Memory Usage (MB) Throughput (q/s)
Classification Dense 0.00% 50030 147.01
ShortGPT/LaCo 35.31% 44624 (0.89x) 221.61 (1.51x)
2SSP 34.77% 44985 (0.90x) 104.03 (0.71x)
SliceGPT 35.16% 42440 (0.85x) 173.94 (1.18x)
Compact 34.99% 32066 (0.64x) 201.19 (1.37x)
Generation Dense 0.00% 21787 81.18
ShortGPT/LaCo 35.31% 16336 (0.75x) 128.03 (1.57x)
Compact 34.99% 14248 (0.65x) 112.40 (1.38x)
Settings: i) Text classification: When running our pruned models on our downstream performance benchmarks, we record the maximum memory usage as well as the throughput, measured in number of questions per second. For this test, we use the HellaSwag benchmark, as it is the longest test in our benchmark suite, which allows us to better discriminate between the methods. For the same reason, we test on LLaMA 3.1-8B using a 3-run average. All models are loaded in 16-bit precision on a single A100-80GB GPU, with a batch size of 256. ii) Text generation: We use the vLLM library (Kwon et al., 2023) to test the memory reduction and inference speedup of our method. As mentioned before, because SliceGPT and 2SSP are incompatible with vLLM, we omit it from our tests. Similarly to the text classification test, we use a 3-run average of LLaMA 3.1-8B in 16-bit precision on a single A100-80GB GPU, but with a batch size of 1 instead, as this is a more realistic workload for text generation. Tests are conducted with 128 input tokens and 128 output tokens.
Our results are in Table 5. Note that ShortGPT and LaCo prune the same number of layers, so the resulting pruned model has an identical architecture. As a result, they have the same inference performance. In the text classification task, our method achieves the highest memory reduction, and the second-highest throughput increase. We attribute the low memory usage to the reduced vocabulary size. During the forward pass, the model logits are stored in GPU memory, which has size $\text{Batch Size}\times\text{Seq Len}\times\text{Vocab Size}$. This matrix can become very large in scenarios with high batch sizes, reaching the tens of gigabytes in this test. By pruning the vocabulary size, Compact significantly reduces the memory usage over other pruning methods. Compact has higher throughput than our width pruning baselines SliceGPT and 2SSP, although it falls behind our depth pruning methods. This mirrors findings in Bian et al. (2025), which showed that scaling down the number of layers results in a proportional increase in throughput, but scaling down the size of each layer does not. This is because matrix multiplications in each layer are performed in parallel, and reducing the matrix size only results in lower GPU utilization with a minimal throughput increase. These results suggest that while Compact achieves faster inference compared to other width pruning methods, further work is needed to match the inference speed of depth pruning methods. In the generation task, the trends in memory usage and throughput are similar to that of the classification task.
4.4 Recovery Fine-tuning
Although Compact is a training-free method, pruned models can be fine-tuned afterwards to recover performance. Due to the page limit, our recovery-finetuning experiments can be found at A.4.
5 Ablation Study
In ablation study, we try to further three questions: i) $\mathcal{Q}_{1}$: how effective our intermediate pruning method Common-act2? ii) $\mathcal{Q}_{2}$: How much the various configuration of Vocab-pruning and FFN-pruning lead to? iii) $\mathcal{Q}_{3}$: how much calibration data is needed?
Answer to $\mathcal{Q}_{1}$: effectiveness of Common-act2.
We compare our novel Common-act2 method with act2. To isolate the effect of the intermediate pruning method, we prune rare vocabulary for all models, then apply the specified intermediate pruning method. We also test another commonly used method, ||act||, which is similar to act2 but uses the summed absolute activations instead of squared activations. Our results are summarized in Figure 2. We find that Common-act2 achieves the highest mean performance compared to our baselines.
Figure 2: Comparison of different intermediate pruning methods on Qwen 2.5-0.5B at a 35% pruning ratio when combined with rare vocabulary removal. We see that common act2 outperforms both act2 and ||act||.
common act2act2||act||66666868707070.470.469.269.267.667.6Avg%
Answer to $\mathcal{Q}_{2}$: vocabulary-intermediate tradeoff
Compact has two hyperparameters: the new vocabulary size $V^{\prime}$ and the new intermediate size $I^{\prime}$. These hyperparameters can be adjusted to produce many different model configurations at a given pruning ratio. We test different configurations of Qwen 2.5-0.5B at a 35% pruning ratio by starting with $V^{\prime}=151936,I^{\prime}=2048$. This configuration leaves $V^{\prime}$ unchanged, which is identical to act2. Then, we progressively increase $I^{\prime}$ and decrease $V^{\prime}$ until $V^{\prime}$ reaches 0. This results in 8 different model configurations to test.
The results can be seen in Table 6. Despite act2’s simplicity, it already achieves better performance than our baselines on its own, without any rare token removal. However, we find that the best result is achieved with a combination of Vocab-pruning and FFN-pruning, validating the effectiveness of our proposed method.
Table 6: Multiple values of $V^{\prime}$ and $I^{\prime}$ over a fixed pruning ratio for Qwen 2.5-0.5B.
V’ I’ Para (M) Ratio Avg Avg% Dense 151936 4864 494 0.00% 50.2 100.0 act2 151936 2048 312 36.84% 31.6 63.0 Compact 131584 2304 311 37.04% 31.8 63.3 111104 2560 309 37.45% 33.1 66.1 90752 2944 315 36.23% 34.2 68.2 70400 3200 314 36.44% 34.5 68.9 49536 3456 311 37.04% 35.3 70.4 29568 3712 310 37.25% 34.7 69.2 9088 3968 308 37.65% 32.4 64.7
Answer to $\mathcal{Q}_{3}$: calibration data size.
Figure 3: Downstream performance across different calibration dataset sizes. Although all benchmarks used 256 calibration samples, 16 samples gives similar performance.
816326412825651210246666686870707272# SamplesAvg%
All of our experiments were conducted using 256 calibration samples from the C4 dataset. We perform ablations over the number of calibration samples used, with our results in Figure 3. We find that Compact is highly robust to the number of calibration samples, and similar performance can be achieved with just 16 samples, implying that the pruning time can be reduced even further.
6 Conclusion
We proposed Compact, a training-free pruning method combining vocabulary and FFN pruning under the common-token distribution. Experiments show that it achieves robust performance across model scales, offering smooth degradation, strong efficiency, and broad deployment compatibility.
Appendix A Appendix
A.1 Parameter Distribution Across Other Model Families
Figures 5 and 5 provide parameter distributions for the LLaMA 3 and Gemma 3 model families, respectively. We see that these models follow the same trend as Qwen 2.5 where smaller LLMs have a higher proportion of vocabulary parameters, corroborating our theoretical analysis.
Figure 4: Llama 3 Parameter Distribution
1B3B8B70B405B00.20.20.40.40.60.60.80.811ProportionVocabFFNAttention
Figure 5: Gemma 3 Parameter Distribution
270M1B4B12B27B00.20.20.40.40.60.60.80.811ProportionVocabFFNAttention
A.2 Evaluation Methodology
Downstream evaluations were conducted using the LM-Evaluation-Harness (Gao et al., 2024a). The evaluation details for each benchmark are in Table 7.
Table 7: Evaluation methodology for each benchmark.
Benchmark n-shot Type Metric MMLU 0 multiple-choice acc HellaSwag 0 multiple-choice acc_norm WinoGrande 0 multiple-choice acc ARC-C 0 multiple-choice acc_norm ARC-E 0 multiple-choice acc_norm PIQA 0 multiple-choice acc_norm GSM8K 5 generative strict_match
A.3 Pruning Hyperparameters
We provide $V^{\prime}$ and $I^{\prime}$ for all our main results in Table 8. These hyperparameters were found by sweeping over all possible configurations for the given pruning ratio, similarly to Table 6.
Table 8: Compact pruning hyperparameters for all main results.
Ratio (%) V’ I’ Qwen 2.5–0.5B 0.00 151936 4864 10.00 99968 4736 20.00 49536 4736 35.00 49536 3456 LLaMA 3.2–1B 0.00 128256 8192 10.00 67968 8192 20.00 56704 7168 35.00 33792 5760 Gemma 3–1B 0.00 262144 6912 10.00 174592 6912 20.00 86912 6912 35.00 95232 5120 LLaMA 3.1–8B 0.00 128256 14336 10.00 73216 13440 20.00 67328 11520 35.00 67840 8448 LLaMA 3.1–70B 0.00 128256 28672 10.00 112384 25216 20.00 111872 21632 35.00 110976 16256
A.4 Recovery Fine-tuning
Although Compact is a training-free method, we can optionally perform recovery fine-tuning after pruning to regain additional performance. To test this, we perform full finetuning on Qwen 2.5-0.5B after pruning at a 35% pruning ratio. We conduct two experiments with this setup:
  1. 1.
    Continued Pretraining (CPT): We use the first 900 million tokens of the FineWeb-Edu dataset (Penedo et al., 2024), a high quality pretraining dataset that achieves SOTA performance compared to other datasets using an equal number of tokens.
  2. 2.
    Self-Data Distillation (SDD): Modern model families like Qwen use very high-quality, private pretraining datasets that have a different distribution than open-source versions. As a result, performing continued pretraining could result in worse performance. We use a modified version of self-data distillation (Thangarasa et al., 2024), which creates a dataset with the same distribution as the training dataset.
    More specifically, we use the unpruned model to generate 900 million tokens of text completions. These completions require at least one input token to begin generation, so we use the first token of examples in FineWeb-Edu. We generate these completions with temperature=1.0 and no other sampling parameters to ensure that the dataset reflects the unmodified probability distribution of the unpruned model. After this, we fine-tune our pruned model on this synthetic dataset.
Our results are shown in Table 9. By manually examining samples from the synthetic dataset, we find that many of the completions are incoherent, often degenerating into repetitive text and hallucinations. This is due to using a tiny teacher model in conjunction with our lack of sampling parameters. Despite this, we find that self-data-distillation significantly outperforms CPT, achieving 5% higher performance recovery and 13% total performance recovery compared to the training-free method. Although CPT achieves higher scores in a majority of the benchmarks, it does not achieve any improvements in MMLU or GSM8K, our two hardest benchmarks. In contrast, SDD is able to consistently improve on all benchmarks, enabling a higher average score. These results highlight that SDD outperforms regular fine-tuning even at the sub-1B parameter range, extending the results in Thangarasa et al. (2024). Notably, our pruned model fine-tuned on 900 million tokens outperforms the recently-released Gemma 3-270M, despite the latter being pretrained on 6 trillion tokens. These results highlight the practicality of pruning techniques compared to pretraining.
Table 9: Comparison of post-pruning recovery fine-tuning methods.
Method Params (M) MMLU HeSw WiGr ARC-C ARC-E PIQA GSM8K Avg Avg% Dense 494 47.3 52.2 56.4 32.3 58.2 69.9 34.9 50.2 100.0 Random - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 57.0 Ours 311 25.5 40.0 53.8 25.2 40.0 62.2 0.5 35.3 70.4 + SDD 311 39.7 43.0 54.9 28.6 50.6 65.6 11.3 41.9 83.6 + CPT 311 25.8 45.1 55.8 27.9 52.7 67.7 0.0 39.3 78.3 Gemma 3-270M 268 26.2 41.5 53.8 28.4 57.2 68.4 1.2 39.5 -